Grade 11Mechanics

a bullet of mass 20g is found to pass two points 30m apart in a time interval of 4 seconds. calculate the kinetic energy of the bullet if it moves with constant speed.

shruti
10 Years agoGrade 11

3 Answers

Yash Arora
10 Years ago
velocity of bullet = 30/4 = 15/2
k.e = 20 *15*15 /(2*2*1000) =1.125 j 
where k.e is kinetic energy
j is joules 
 
velocity = distance / time 
Sanchita
9 Years ago
Mass=0.2 kgV= dist \timeSo velocity = 15/2 m/sKE=1/2 m v.v 1/2 *0.2*15/2*15/20.05625 JWhere ke is kinetic energyJ is in joules and mass in kgs
Sd
8 Years ago
Mass is twenty grams . Two points separated by a distance thirty meters.time is four seconds ke is half m v square